In 1920, a 6.5 × 104 kg meteorite was found in Africa. Suppose a meteorite with this mass enters Earth’s atmosphere with a speed of 1.0 km/s. What is the change in the meteorite’s momentum if an average constant force of −1.7 × 106 N acts on the meteorite for 30.0 s? How far does the meteorite travel during this time? 


1
Expert's answer
2022-01-10T09:09:34-0500

"\Delta p=mv_1-Ft=14\cdot 10^6~\frac{kg\cdot m}s,"

"v_2=\frac{\Delta p}m=215~\frac ms,"

"s=\frac{v_1+v_2}2t=18.2~km."